OpenStudy (anonymous):

-1.77846154 as a fraction

OpenStudy (anonymous):

use calculator?!

OpenStudy (anonymous):

if you have a graphing calculator

OpenStudy (anonymous):

I have one, I just have no idea where it is...

OpenStudy (anonymous):

ahh alright hold on let me grab mine

OpenStudy (anonymous):

Thanks! :D

OpenStudy (anonymous):

it is not a fraction

OpenStudy (anonymous):

88923077/50000000

OpenStudy (anonymous):

I think that's in reduced form...

OpenStudy (anonymous):

so, basically, the problem is: y-4.36 =(2.89/13) (x-8)

OpenStudy (anonymous):

oops I meant -88923077/50000000

OpenStudy (anonymous):

the (2.89/13) is the fraction. So am i just going about this wrong?

OpenStudy (anonymous):

what are you trying to do with the problem?

OpenStudy (anonymous):

2.89/13 is not 1.7ish... it's much more less

OpenStudy (anonymous):

2.89/13=289/1300

OpenStudy (anonymous):

what are you trying to find? the line?

OpenStudy (anonymous):

\[y=\frac{289}{1300}x+\frac{839}{325}\]

OpenStudy (anonymous):

I figured it out, nevermind, but thanks anyways. I was trying to get this: y-7.25=0.17(x-25) y-7.25=0.17x-4.25 y=0.17x+3